# Cascading from the winds to the disk: the universality of supernovae-driven turbulence in different galactic interstellar media

### v2

@@ 3.4 Interstellar medium phases
Figure 7: One-dimensional, isotropic power spectra normalized by the integrated power, as a function of wavenumber, $k$, normalized by the system scale wavenumber, $2\pi/L$, with the gaseous scale height, $k_{0}\sim\ell_{0}^{-1}$, annotated on each panel with the grey vertical line. The shaded regions show the $1\sigma$ fluctuations of the spectra in time. We show MW (left column) and SB (right column), for both seeding mechanisms (dark and light colors). All fitted indexes for spectra are shown in Table 3. Row one: the velocity spectrum compensated by $k^{-5/3}$ for Kolmogorov (1941)-type turbulence. Row two: the velocity spectrum decomposed into compressible $\bm{u}_{c}$ (dashed) and incompressible $\bm{u}_{s}$ (solid) modes for each seeding mechanism (offset for clarity), compensated by $k^{-2}$ Burgers (1948)-type turbulence, and with annotations for the quasi-acoustic turbulence $k^{-3}$ (Wang et al., 2017; Beattie et al., 2025b). Row three: the kinetic energy spectrum compensated by $k^{-4/3}$ (found previously in compressible turbulence Grete et al. 2021). Row four: the $c_{s}$ spectrum, with $\bm{u}_{c}$ velocity spectra modes for each simulation plotted in dashed black and grey lines, showing that both spectra trace one another other (a natural repercussion of being in the adiabatic limit and mediated by acoustic waves, Section 4.3), following $\sim k^{-2}$ for the MW simulations. We show further relevant $k$-scaling annotations in black for each panel, which we discuss in the main text.
4 Power spectra and turbulence in the galaxy models
@@ 4.1 Definitions
4.2 The $k$-space structure of the turbulence
@@
In the third row we plot the kinetic energy spectra compensated by $k^{-4/3}$, found previously in compressible (magnetized) turbulence (Grete et al., 2021). For $k\lesssim\ell_{0}^{-1}$, we see close alignment between the $k^{-4/3}$ model and $d\rho u^{2}(k)/dk$. For SB, we see this behavior at small and large scales, but with a transition between at $k\approx\ell_{0}^{-1}$. This transition is most likely due to the gaseous profile of the disk imprinting itself across the spectrum. As discussed in the Appendix of Beattie et al. (2025c), this increase in kinetic energy density at high $k$, compared to $du^{2}(k)/dk$, is likely due to high-density clumps on small scales. The velocity power spectrum is more universal across scales and between models, suggesting the velocity modes are correlated through the turbulence in the winds and disk, which we will explore in more detail in Section 4.4.
4.3 What sets the large-scale phase structure in the ISM?
@@
4.4 Cascading from the winds to the disk
@@
We find that all spectra collapse across galactic models and seeding schemes into just a single velocity (top) and kinetic (bottom) spectrum, that are smooth across $\ell^{-1}_{0}$. The $du^{2}(k)/dk\propto k^{-3/2}$ discussed in Section 4.2 and previously measured in Padoan et al. (2016) and Beattie et al. (2025c) extends across both simulations, connecting the winds to the turbulent disk in a continuous manner. For the $d\rho u^{2}(k)/dk$ spectrum we find a broken power law, that seems to break at roughly $k\ell_{0}/2\pi\approx 1$, with $d\rho u^{2}(k)/dk\propto k^{-1/2}$ on $k\ell_{0}/2\pi\lesssim 1$ and $d\rho u^{2}(k)/dk\propto k^{-4/3}$ on $k\ell_{0}/2\pi\gtrsim 1$. This is likely due to the gaseous disk imprinting itself on the power spectrum, due to the mass density contribution in the spectrum. Federrath (2013) previously found that the $d\rho u^{2}(k)/dk$ was sensitive to the underlying driving field and $\mathcal{M}$, whereas the $du^{2}(k)/dk$ was not.
@@ 5 Summary and Conclusions
  • In Section 3 we show the notably different dynamical, thermal and density structure in the ISMs between models. The SB models are wind-dominated, characterized by a thin gaseous scale height ($\ell_{0}\approx 40\,{\rm{pc}}$; which we model with an isothermal, pressure-balanced atmosphere model; Equation 14), a prominent hot galactic wind with large thermal fluctuations, and a short turbulent turnover time on $\ell_{0}$ of order $t_{0}\approx 0.3\,\rm{Myr}$. The MW model is, on the other hand, disk-dominated with a thick disk ($\ell_{0}\approx 120$pc), a very tenuous wind with lower thermal and velocity fluctuations, and a longer turbulent turnover time on $\ell_{0}$ of order $t_{0}\approx 4\,\rm{Myr}$. The FX seeding models lead to a more dynamic, almost adiabatic wind, with larger amplitude thermal and turbulent velocity fluctuations compared to the SC models.
@@ Acknowledgments
Figure 9: Vertical disk profile fits (light purple) for each model (top left annotation; dark purple; also shown in Figure 3) using the isothermal pressure-equilibrium profile model (Equation 17), with free parameter $\sigma_{z}^{2}$. We use this disk-component model to calculate gaseous scale height $\ell_{0}$ (Equation 18) detailed in Section 3.2. To fit, we use maximum likelihood fitting, constructing the posterior by utilizing emcee (Foreman-Mackey et al., 2013), method detailed in Appendix A.
@@ Appendix A Error propagation and fitting the galactic profile models
We model the vertical density profile of the disk using Equation 17 (detailed in Section 3.2) and fit the free parameter $\sigma_{z}^{2}$ (Equation 16). We employ a MCMC maximum-likelihood method using the emcee (Foreman-Mackey et al., 2013) ensemble sampler to construct the posterior. The likelihood was constructed from the log-density profiles $\log\rho(z)$, weighted by the variance at each height $z$, and an additional Gaussian window function to extract only the disk component. We pick uniform priors: $10^{4}\,\rm{cms}^{-1}<\sigma_{z}<8\times 10^{6}\,\rm{cms}^{-1}$, based on physical constraints for the turbulent and thermally supported disk. For each model we ran $50$ chains for 5,0005,\!000 steps, discarding the first $20\%$ of the chains as burn-in stage. We report the median and standard deviation of $\sigma_{z}$ as the best-fit value and uncertainty directly from the posterior. From $\sigma_{z}$ and Equation 18, we derive the gaseous scale height $\ell_{0}$, and propagate the uncertainty by applying standard Gaussian error propagation, where
@@ Appendix B Fitted Power Law Indexes For Spectra
Note. — Column (1): Numerical model label, where MW means Milky Way analogue, SB means starburst analogue, SC refers to the self-consistent SNe seeding prescription, and FX refers to the fixed seeding prescription (see Section 2.2.2 for more details about the exact seeding prescription). All indexes are derived from fits across the range of modes $kL/2\pi\in[8.5,25]$. Column (2): The velocity power spectrum. Column (3): The incompressible velocity mode power spectrum. Column (4): The compressible velocity mode power spectrum. Column (5): The kinetic energy power spectrum. Column (6): The sound speed power spectrum.
In Table 3 we show the power law indexes derived by fitting to the scaling ranges of each of the spectra discussed in Section 4 and plotted in Figure 7. For the fits, we use least-squares in $\log-\log$, propagating the $1\sigma$ from the temporal averaging for each of the spectra. As shown in Beattie et al. (2025c), it is not obvious that there is a true inertial (constant $u^{3}/\ell$) range of modes, so we therefore just pick the range $kL/2\pi\in[8.5,25]$ by eye, and perform the fit over that range. Because these simulations are not scale-free, i.e., the box is stratified with scale heights in all quantities (as we show in Figure 3 and Figure 9), we are unable to directly compare the indexes between the galaxy models because $kL/2\pi\in[8.5,25]$ is probing different physical scales in each of the simulations (see Figure 8 for how to properly combine and compare the spectra across galaxies). However, for the same galactic models, with different seeds, we see broad agreement within $1\sigma$, as expected qualitatively from Figure 7.
@@ Appendix C Divergence and Compressible Mode Power Spectrum
In Section 4 we show the compressible mode $\bm{u}_{c}$ spectrum was directly related to the sound speed spectrum, which was controlling the ISM phase structure in our models with a thick disk (Milky Way analogue models). Here we formalize the connection with $\bm{u}_{c}$ and the fluid divergence. The divergence $\bm{\nabla}\cdot\bm{u}$ and $\bm{u}_{c}$ are intrinsically related because only $\bm{u}_{c}$ satisfies $\bm{\nabla}\cdot\bm{u}\neq 0$. In Figure 10 we plot the power spectrum of both quantities (left MW models, right SB models), with the expectation that $du_{c}^{2}(k)/dk\propto d(\bm{\nabla}\cdot\bm{u})^{2}(k)/(k^{2}dk)$ based on dimensionality. Indeed, we find that $du_{c}^{2}(k)/dk\propto d(\bm{\nabla}\cdot\bm{u})^{2}(k)/(k^{2}dk)$ across all resolved modes in the simulations. As we show in Figure 7, $du_{c}^{2}(k)/dk\propto k^{-2}$, which we associate with Burgers (1948) turbulence (as in Beattie et al. 2025a on large scales).
Figure 10: One-dimensional, isotropic velocity divergence, $\bm{\nabla}\cdot\bm{u}$, power spectrum (blues), and compressible mode, $\bm{u}_{c}$, power spectrum (dashed lines) normalized by the integrated power, for each seeding mechanism, for MW (left) and SB (right) galaxy models. We presented in Section 4.3 and Figure 7 (fourth row) that the sound speed scales with the compressible mode spectrum, explaining the temperature fluctuations in the medium. Further, as expected, we show that the divergence spectrum and the compressible mode spectrum are simply related by a $k^{2}$ scaling to get the divergence into the correct dimensions.
@@ Appendix D Convergence test
For the purpose of testing the convergence of our spectral results, we calculate the velocity power spectra at each resolution ($N_{\rm grid}^{3}=128^{3},\;256^{3},\;\text{and}\;512^{3}$) for all models. We show this in Figure 11, with MW models on the left, and SB models on the right, and SC and FX runs shown in purples and pinks, respectively. We normalize by the integral power and compensate by $k^{-5/3}$, as in Section 4. It is visible that as we increase the resolution, we resolve an extended range of self-similar modes in the spectrum, increasing the power at higher $k$ as the numerical dissipation scale moves to smaller scales. For implicit large-eddy simulations, this is exactly the behavior we expect because the viscous truncation of the cascade is purely a function of the grid (Grete et al., 2023; Shivakumar & Federrath, 2025; Grehan et al., 2025). We find $128^{3}$ is insufficient to resolve any of the self-similar range of modes, but as we go to higher resolutions, we find that the self-similar range of modes emerges, and at $512^{3}$ we have almost an order of magnitude of modes in this self-similar range. Furthermore, the spectra begin to roughly follow $\sim k^{-3/2}$. The key point is that the spectra begin to converge to a similar shape as we go to higher resolution, which means that the measurements we make in this study are capturing some of the asymptotic dynamics of the SNe-driven turbulence.

### v3

@@ III.4 Interstellar medium phases
Figure 7: One-dimensional, isotropic power spectra normalized by the integrated power, as a function of wavenumber, $k$, normalized by the system scale wavenumber, $2\pi/L$, with the gaseous scale height, $k_{0}\sim\ell_{0}^{-1}$, annotated on each panel with the grey vertical line. The shaded regions show the $1\sigma$ fluctuations of the spectra in time. We show MW (left column) and SB (right column), for both seeding mechanisms (dark and light colors). All fitted indexes for spectra are shown in Table 3. Row one: the velocity spectrum compensated by $k^{-5/3}$ for Kolmogorov (1941)-type turbulence. Row two: the velocity spectrum decomposed into compressible $\bm{u}_{c}$ (dashed) and incompressible $\bm{u}_{s}$ (solid) modes for each seeding mechanism (offset for clarity), compensated by $k^{-2}$ Burgers (1948)-type turbulence, and with annotations for the quasi-acoustic turbulence $k^{-3}$ (Wang et al., 2017; Beattie et al., 2025b). Row three: the kinetic energy spectrum compensated by $k^{-4/3}$ (found previously in compressible turbulence Grete et al. 2021). Row four: the $c_{s}$ spectrum, with $\bm{u}_{c}$ velocity spectra modes for each simulation plotted in dashed black and grey lines, showing that both spectra trace one another other (a natural repercussion of being in the adiabatic limit and mediated by acoustic waves, Section IV.3), following $\sim k^{-2}$ for the MW simulations. We show further relevant $k$-scaling annotations in black for each panel, which we discuss in the main text.
IV Power spectra and turbulence in the galaxy models
@@ IV.1 Definitions
IV.2 The $k$-space structure of the turbulence
@@
In the third row we plot the kinetic energy spectra compensated by $k^{-4/3}$, found previously in compressible (magnetized) turbulence (Grete et al., 2021). For $k\lesssim\ell_{0}^{-1}$, we see close alignment between the $k^{-4/3}$ model and $d\rho u^{2}(k)/dk$. For SB, we see this behavior at small and large scales, but with a transition between at $k\approx\ell_{0}^{-1}$. This transition is most likely due to the gaseous profile of the disk imprinting itself across the spectrum. As discussed in the Appendix of Beattie et al. (2025c), this increase in kinetic energy density at high $k$, compared to $du^{2}(k)/dk$, is likely due to high-density clumps on small scales. The velocity power spectrum is more universal across scales and between models, suggesting the velocity modes are correlated through the turbulence in the winds and disk, which we will explore in more detail in Section IV.4.
IV.3 What sets the large-scale phase structure in the ISM?
@@
IV.4 Cascading from the winds to the disk
@@
We find that all spectra collapse across galactic models and seeding schemes into just a single velocity (top) and kinetic (bottom) spectrum, that are smooth across $\ell^{-1}_{0}$. The $du^{2}(k)/dk\propto k^{-3/2}$ discussed in Section IV.2 and previously measured in Padoan et al. (2016) and Beattie et al. (2025c) extends across both simulations, connecting the winds to the turbulent disk in a continuous manner. For the $d\rho u^{2}(k)/dk$ spectrum we find a broken power law, that seems to break at roughly $k\ell_{0}/2\pi\approx 1$, with $d\rho u^{2}(k)/dk\propto k^{-1/2}$ on $k\ell_{0}/2\pi\lesssim 1$ and $d\rho u^{2}(k)/dk\propto k^{-4/3}$ on $k\ell_{0}/2\pi\gtrsim 1$. This is likely due to the gaseous disk imprinting itself on the power spectrum, due to the mass density contribution in the spectrum. Federrath (2013) previously found that the $d\rho u^{2}(k)/dk$ was sensitive to the underlying driving field and $\mathcal{M}$, whereas the $du^{2}(k)/dk$ was not.
@@ V Summary and Conclusions
  • In Section III we show the notably different dynamical, thermal and density structure in the ISMs between models. The SB models are wind-dominated, characterized by a thin gaseous scale height ($\ell_{0}\approx 40\,{\rm{pc}}$; which we model with an isothermal, pressure-balanced atmosphere model; Equation 14), a prominent hot galactic wind with large thermal fluctuations, and a short turbulent turnover time on $\ell_{0}$ of order $t_{0}\approx 0.3\,\rm{Myr}$. The MW model is, on the other hand, disk-dominated with a thick disk ($\ell_{0}\approx 120$pc), a very tenuous wind with lower thermal and velocity fluctuations, and a longer turbulent turnover time on $\ell_{0}$ of order $t_{0}\approx 4\,\rm{Myr}$. The FX seeding models lead to a more dynamic, almost adiabatic wind, with larger amplitude thermal and turbulent velocity fluctuations compared to the SC models.
@@ Acknowledgments
Figure 9: Vertical disk profile fits (light purple) for each model (top left annotation; dark purple; also shown in Figure 3) using the isothermal pressure-equilibrium profile model (Equation 17), with free parameter $\sigma_{z}^{2}$. We use this disk-component model to calculate gaseous scale height $\ell_{0}$ (Equation 18) detailed in Section III.2. To fit, we use maximum likelihood fitting, constructing the posterior by utilizing emcee (Foreman-Mackey et al., 2013), method detailed in Appendix A.
@@ Appendix A Error propagation and fitting the galactic profile models
We model the vertical density profile of the disk using Equation 17 (detailed in Section III.2) and fit the free parameter $\sigma_{z}^{2}$ (Equation 16). We employ a MCMC maximum-likelihood method using the emcee (Foreman-Mackey et al., 2013) ensemble sampler to construct the posterior. The likelihood was constructed from the log-density profiles $\log\rho(z)$, weighted by the variance at each height $z$, and an additional Gaussian window function to extract only the disk component. We pick uniform priors: $10^{4}\,\rm{cms}^{-1}<\sigma_{z}<8\times 10^{6}\,\rm{cms}^{-1}$, based on physical constraints for the turbulent and thermally supported disk. For each model we ran $50$ chains for 5,0005,\!000 steps, discarding the first $20\%$ of the chains as burn-in stage. We report the median and standard deviation of $\sigma_{z}$ as the best-fit value and uncertainty directly from the posterior. From $\sigma_{z}$ and Equation 18, we derive the gaseous scale height $\ell_{0}$, and propagate the uncertainty by applying standard Gaussian error propagation, where
@@ Appendix B Fitted Power Law Indexes For Spectra
Note. — Column (1): Numerical model label, where MW means Milky Way analogue, SB means starburst analogue, SC refers to the self-consistent SNe seeding prescription, and FX refers to the fixed seeding prescription (see Section II.2.2 for more details about the exact seeding prescription). All indexes are derived from fits across the range of modes $kL/2\pi\in[8.5,25]$. Column (2): The velocity power spectrum. Column (3): The incompressible velocity mode power spectrum. Column (4): The compressible velocity mode power spectrum. Column (5): The kinetic energy power spectrum. Column (6): The sound speed power spectrum.
In Table 3 we show the power law indexes derived by fitting to the scaling ranges of each of the spectra discussed in Section IV and plotted in Figure 7. For the fits, we use least-squares in $\log-\log$, propagating the $1\sigma$ from the temporal averaging for each of the spectra. As shown in Beattie et al. (2025c), it is not obvious that there is a true inertial (constant $u^{3}/\ell$) range of modes, so we therefore just pick the range $kL/2\pi\in[8.5,25]$ by eye, and perform the fit over that range. Because these simulations are not scale-free, i.e., the box is stratified with scale heights in all quantities (as we show in Figure 3 and Figure 9), we are unable to directly compare the indexes between the galaxy models because $kL/2\pi\in[8.5,25]$ is probing different physical scales in each of the simulations (see Figure 8 for how to properly combine and compare the spectra across galaxies). However, for the same galactic models, with different seeds, we see broad agreement within $1\sigma$, as expected qualitatively from Figure 7.
@@ Appendix C Divergence and Compressible Mode Power Spectrum
In Section IV we show the compressible mode $\bm{u}_{c}$ spectrum was directly related to the sound speed spectrum, which was controlling the ISM phase structure in our models with a thick disk (Milky Way analogue models). Here we formalize the connection with $\bm{u}_{c}$ and the fluid divergence. The divergence $\bm{\nabla}\cdot\bm{u}$ and $\bm{u}_{c}$ are intrinsically related because only $\bm{u}_{c}$ satisfies $\bm{\nabla}\cdot\bm{u}\neq 0$. In Figure 10 we plot the power spectrum of both quantities (left MW models, right SB models), with the expectation that $du_{c}^{2}(k)/dk\propto d(\bm{\nabla}\cdot\bm{u})^{2}(k)/(k^{2}dk)$ based on dimensionality. Indeed, we find that $du_{c}^{2}(k)/dk\propto d(\bm{\nabla}\cdot\bm{u})^{2}(k)/(k^{2}dk)$ across all resolved modes in the simulations. As we show in Figure 7, $du_{c}^{2}(k)/dk\propto k^{-2}$, which we associate with Burgers (1948) turbulence (as in Beattie et al. 2025a on large scales).
Figure 10: One-dimensional, isotropic velocity divergence, $\bm{\nabla}\cdot\bm{u}$, power spectrum (blues), and compressible mode, $\bm{u}_{c}$, power spectrum (dashed lines) normalized by the integrated power, for each seeding mechanism, for MW (left) and SB (right) galaxy models. We presented in Section IV.3 and Figure 7 (fourth row) that the sound speed scales with the compressible mode spectrum, explaining the temperature fluctuations in the medium. Further, as expected, we show that the divergence spectrum and the compressible mode spectrum are simply related by a $k^{2}$ scaling to get the divergence into the correct dimensions.
@@ Appendix D Convergence test
For the purpose of testing the convergence of our spectral results, we calculate the velocity power spectra at each resolution ($N_{\rm grid}^{3}=128^{3},\;256^{3},\;\text{and}\;512^{3}$) for all models. We show this in Figure 11, with MW models on the left, and SB models on the right, and SC and FX runs shown in purples and pinks, respectively. We normalize by the integral power and compensate by $k^{-5/3}$, as in Section IV. It is visible that as we increase the resolution, we resolve an extended range of self-similar modes in the spectrum, increasing the power at higher $k$ as the numerical dissipation scale moves to smaller scales. For implicit large-eddy simulations, this is exactly the behavior we expect because the viscous truncation of the cascade is purely a function of the grid (Grete et al., 2023; Shivakumar & Federrath, 2025; Grehan et al., 2025). We find $128^{3}$ is insufficient to resolve any of the self-similar range of modes, but as we go to higher resolutions, we find that the self-similar range of modes emerges, and at $512^{3}$ we have almost an order of magnitude of modes in this self-similar range. Furthermore, the spectra begin to roughly follow $\sim k^{-3/2}$. The key point is that the spectra begin to converge to a similar shape as we go to higher resolution, which means that the measurements we make in this study are capturing some of the asymptotic dynamics of the SNe-driven turbulence.